# Directed-Tokens: A Robust Multi-Modality Alignment Approach to Large Language-Vision Models

### v1

@@ 1 Introduction
Large multimodal models (LMMs) have gained more attention recently due to their outstanding capabilities in general-purpose assistants. By training via visual instruction tuning [39, 37, 38, 3, 29, 28], these LMMs, e.g., LLaVA [39], InstructBLIP [29, 28], MiniGPT-4 [69], Qwen-VL [4], have shown impressive performance on instruction-following and visual reasoning tasks. Then, these LMMs have been further developed for specific scientific tasks, e.g., LLaVA-Med [26], DriveGPT4 [56], PMC-LLaMA [53], etc. To measure the performance and capabilities of LMMs, several benchmarks [63, 58] have been introduced to evaluate the LMMs in different aspects, e.g., art, business, health and medicine, science, tech and engineering, and other fields. Recent studies further improve the performance of LMMs by scaling the training data [37, 38, 66, 65], using better visual encoders [5, 28, 28, 51] or large language models [25], improving objective learning [50, 31, 60, 68], using multimodal preference data [54, 45, 61], extending to other modalities [35, 34, 33] (e.g., videos [35, 34, 43], graphs [33]).
@@ 2 Related Work
Large Multimodal Models. Thanks to the remarkable advancements in LLMs [34, 10, 1, 46, 3], it has promoted the development of LMMs. Numerous LMMs have been developed to enhance effectiveness in handling such data, which can be classified into large vision-language models [39, 37], large video-language models [52, 67, 35, 30], and large audio-language models [20, 15]. Early work by [2] effectively bridged vision and language modalities in a few-shot learning setting, followed by [28], which enhanced inter-modality connectivity via Q-Former. This was further developed into an instruction-aware model within the vision-language instruction-tuning framework [39]. LLaVA [39] established a streamlined visual-to-language space projection using a linear layer, later refined by [37] with an MLP and AnyRes, a technique adept at handling high-resolution images. Subsequent studies [38, 24, 64, 23, 27] contributed further improvements, culminating in a robust model [25] capable of handling diverse vision tasks, including video comprehension. Later, [26] extended LLaVA to a biomedical model through a cost-effective curriculum learning approach. [7] leveraged multimodal federated learning to enhance LMM training. Meanwhile, [57, 55, 36] introduced LMMs for 3D point cloud understanding tasks.
Shuffle Learning. Decomposing an image into smaller patches, shuffling these patches, and then reconstructing their original order forms a classical pattern recognition problem known as the jigsaw puzzle. The early work in classification [14] solved the jigsaw puzzle by predicting the spatial position of each patch. Later, it has proven highly useful in self-supervised learning for feature representation. Indeed, [6] incorporates a jigsaw puzzle challenge alongside classification to improve visual information generalization across domains. [12] employs a jigsaw puzzle generator to create varying levels of granularity for fine-grained classification. Chen et al. [9] extended shuffle learning to transformers to enhance performance in visual recognition tasks. Truong et al. [48] developed a robust video understanding model by solving the shuffled temporal frames via the directed attention mechanism. This body of work suggests that using the jigsaw puzzle as a pretext task is broadly effective for generalizing spatial information within images. Building on this, we investigate the impact of reconstructing the order of elements not only in images but also in the texts, analyzing the interplay between these two modalities in relation to each other’s permutated version.
@@ 3 The Proposed Directed Token Approach to LMM
The success of the current state-of-the-art LMMs [39, 37, 38] relies on auto-regressive modeling. Indeed, the form of auto-regression naturally matches the nature of the data, where each input token in the sequence depends on the previous ones. This modeling approach can capture the complex dependencies and correlations within the image and language and maintain consistency and coherence in data. As a result, the order of data, i.e., visual image patches or textual sentences, plays an important role since the LMM auto-regressively models multimodal inputs conditioning on all previous elements to maintain context and coherence. If the data order is incorrect, the LMM loses its ability to model correct contexts and produce logically consistent predictions. Under this modeling principle, we aim to improve the LMMs by addressing two fundamental questions. First, given a random shuffled image, we aim to develop an LMM capable of reconstructing the original image that is represented by the current textual descriptions (Fig. 3(A)). This learning mechanism encourages the LMM to learn the strong correlation between visual and textual features, thus providing a better understanding of visual information. Second, given an original image and a shuffled textual description, our LMM aims to predict the correct textual description that represents the visual information of the image (Fig. 3(B)). The learning objective allows the model to improve correlation across modalities further, thus enhancing the important role of visual information in the LMM’s responses. In the next sections, we will describe our approach to developing these two learning objectives in the pre-training and fine-tuning phases of the LMMs.

### v2

@@ 1 Introduction
Large multimodal models (LMMs) have gained more attention recently due to their outstanding capabilities in general-purpose assistants. By training via visual instruction tuning [40, 38, 39, 3, 30, 29, 43, 57], these LMMs, e.g., LLaVA [40], InstructBLIP [30, 29], MiniGPT-4 [79], Qwen-VL [4], have shown impressive performance on instruction-following and visual reasoning tasks. Then, these LMMs have been further developed for specific scientific tasks, e.g., LLaVA-Med [27], DriveGPT4 [66], PMC-LLaMA [62], etc. To measure the performance and capabilities of LMMs, several benchmarks [73, 68] have been introduced to evaluate the LMMs in different aspects, e.g., art, business, health and medicine, science, tech and engineering, and other fields. Recent studies further improve the performance of LMMs by scaling the training data [38, 39, 76, 75], using better visual encoders [5, 29, 29, 60] or large language models [26], improving objective learning [59, 32, 70, 78], using multimodal preference data [64, 52, 71], extending to other modalities [36, 35, 34] (e.g., videos [36, 35, 46], graphs [34, 46, 48, 47], human brain signals [50, 63]).
@@ 2 Related Work
Large Multimodal Models. Thanks to the remarkable advancements in LLMs [35, 10, 1, 53, 3, 58, 45], it has promoted the development of LMMs. Numerous LMMs have been developed to enhance effectiveness in handling such data, which can be classified into large vision-language models [40, 38], large video-language models [61, 77, 36, 31], and large audio-language models [21, 16]. Early work by [2] effectively bridged vision and language modalities in a few-shot learning setting, followed by [29], which enhanced inter-modality connectivity via Q-Former. This was further developed into an instruction-aware model within the vision-language instruction-tuning framework [40]. LLaVA [40] established a streamlined visual-to-language space projection using a linear layer, later refined by [38] with an MLP and AnyRes, a technique adept at handling high-resolution images. Subsequent studies [39, 25, 74, 24, 28] contributed further improvements, culminating in a robust model [26] capable of handling diverse vision tasks, including video comprehension. Later, [27] extended LLaVA to a biomedical model through a cost-effective curriculum learning approach. [7] leveraged multimodal federated learning to enhance LMM training. Meanwhile, [67, 65, 37] introduced LMMs for 3D point cloud understanding tasks.
Shuffle Learning. Decomposing an image into smaller patches [12, 55, 56, 49], shuffling these patches [55, 15], and then reconstructing their original order forms a classical pattern recognition problem known as the jigsaw puzzle. The early work in classification [15] solved the jigsaw puzzle by predicting the spatial position of each patch. Later, it has proven highly useful in self-supervised learning for feature representation. Indeed, [6] incorporates a jigsaw puzzle challenge alongside classification to improve visual information generalization across domains. [13] employs a jigsaw puzzle generator to create varying levels of granularity for fine-grained classification. Chen et al. [9] extended shuffle learning to transformers to enhance performance in visual recognition tasks. Truong et al. [55] developed a robust video understanding model by solving the shuffled temporal frames via the directed attention mechanism. This body of work suggests that using the jigsaw puzzle as a pretext task is broadly effective for generalizing spatial information within images. Building on this, we investigate the impact of reconstructing the order of elements not only in images but also in the texts, analyzing the interplay between these two modalities in relation to each other’s permutated version.
@@ 3 The Proposed Directed Token Approach to LMM
The success of the current state-of-the-art LMMs relies on auto-regressive modeling [40, 38, 39, 44]. Indeed, the form of auto-regression naturally matches the nature of the data, where each input token in the sequence depends on the previous ones. This modeling approach can capture the complex dependencies and correlations within the image and language and maintain consistency and coherence in data. As a result, the order of data, i.e., visual image patches or textual sentences, plays an important role since the LMM auto-regressively models multimodal inputs conditioning on all previous elements to maintain context and coherence. If the data order is incorrect, the LMM loses its ability to model correct contexts and produce logically consistent predictions. Under this modeling principle, we aim to improve the LMMs by addressing two fundamental questions. First, given a random shuffled image, we aim to develop an LMM capable of reconstructing the original image that is represented by the current textual descriptions (Fig. 3(A)). This learning mechanism encourages the LMM to learn the strong correlation between visual and textual features, thus providing a better understanding of visual information. Second, given an original image and a shuffled textual description, our LMM aims to predict the correct textual description that represents the visual information of the image (Fig. 3(B)). The learning objective allows the model to improve correlation across modalities further, thus enhancing the important role of visual information in the LMM’s responses. In the next sections, we will describe our approach to developing these two learning objectives in the pre-training and fine-tuning phases of the LMMs.
@@ 5 Conclusions
Acknowledgment. This work is partly supported by NSF CAREER (No. 2442295), NSF SCH (No. 2501021), NSF E-RISE (No. 2445877), NSF SBIR Phase 2 (No. 2247237) and USDA/NIFA Award. We also acknowledge the Arkansas High-Performance Computing Center (HPC) for GPU servers.
Appendices